Timberly Construction negotiates a lump-sum purchase of several assets from a company that is going out of business. The purchase is completed on January 1, 2015, at a total cash price of $800,000 for a building, land, land improvements, and four vehicles. The estimated market values of the assets are building, $472,800; land, $315,200; land improvements, $68,950; and four vehicles, $128,050. The company’s fiscal year ends on December 31.

1.Prepare the journal entry to record the purchase.

2.

Compute the depreciation expense for year 2015 on the building using the straight-line method, assuming a 15-year life and a $30,000 salvage value. (Round your answers to the nearest whole dollar.)

3.

Compute the depreciation expense for year 2015 on the land improvements assuming a five-year life and double-declining-balance depreciation.
Answer #1

1) Journal entry

Date account and explanation Debit Credit
Building (800000*48%) 384000
Land 256000
Land improvement 56000
Vehicle 104000
Cash 800000
(To record lump sum purchase)

2) Depreciation for building = (384000-30000/15) = 23600

3) Depreciation on land improvement = 56000*40% = 22400